OpenStudy (anonymous):

Which of the following is equivalent to |x − 1| = 20?

OpenStudy (anonymous):

x − 1 = 20 x − 1 = 20 or x − 1 = -20 x − 1 = -20 none of the above

OpenStudy (jiteshmeghwal9):

(b)